OpenStudy (anonymous):

[2.06] Solve for x: 2 over 3 (x − 4) = 2x. 2 −2 −8 −4

OpenStudy (anonymous):

2x(3x-12)=2 6x^2-24x-2=0 this is a quadratic equation divide by 2 3x^2-6x-1=0

OpenStudy (anonymous):

can u solve it further ?

OpenStudy (anonymous):

uhhhh no

OpenStudy (anonymous):

use quad formula

OpenStudy (anonymous):

or sum and product method

OpenStudy (anonymous):

3x^2-12x-1=0